OpenStudy (anonymous):

My answer is C Solve. 2x2 + x = 15. (1 points) Answer x = 5/2 and x = -3 x = 2 and x = -15 x = -2 and x = 5/3 x = 17 and x = 13

OpenStudy (anonymous):

2x2 + x = 15 2x2 + x -15 =0 (2x-5)(x+3)=0 x= 5/2 OR x=-3

OpenStudy (anonymous):

You can also use formula, which is much easier if you ask me \[\frac{ -b \pm \sqrt{b ^{2}-4*a*c} }{ 2*a }\] a=2, b=1, c=-15

OpenStudy (anonymous):

aww you're so kind, thank you

OpenStudy (anonymous):

x1 = 2.5 x2 = -3 As @julliemyers said :) Good luck!

OpenStudy (anonymous):

thanks!